solve using the elimination method.
10x + 11y = -26.5
20x + 5y = -27.5

Answer by mananth(16946)

10 x + 11 y = 26.5 .............1
20 x + 5 y = -27.5 .............2

multiply (1)by -5
Multiply (2) by 11
-50 x -55 y = -132.5
220 x 55 y = -302.5
Add the two equations
170 x = -435
/ 170
x = -2.56
plug value of x in (1)
10 x + 11 y = 26.5
-25.59 + 11 y = 26.5
11 y = 26.5 + 25.59
11 y = 52.09
y = 4.74